INAF LBC+MODS queue observing, 2012 February 21-22
 
Observer:  K. Boutsia, M. Fumana 
Telescope Operator: D. Gonzales-Huerta 
Support Astronomer: D. Thompson 
Instrument: LBC+MODS 
 Summary 
Mostly clear skies at sunset. During the whole night the seeing has been variable between 1.3" and 2.0". 
Today we finally managed to obtain the NS data. The PI has been informed and the data are beeing processed.
We tried to observe the 
IC10VI target (german request) but the seeing was >2.0" and completely out of the requested specs and we aborted.
Observed all Donati targets, since the seeing was poor and he had loose constraints. 3 out of 5 of the Donati targets are within specs.
After slewing to Gilli, to find out that also here the first data were out of spec, we aborted and switched to MODS. No standard stars obtained with LBC tonight.
Using MODS, we started with Fiore (Cosmos2) but 1.5hr out of 3.5hr are out of specs (see Fig at the end of log). Standard calibrations obtained. For flat calibrations see repository 20120220UT
At the end of the night we started Donnarumma, which is out of specs also (see Fig at the end of log).
The Donnarumma calibrations for mask1 have as PROPID sdss1029 even though the actual target is A1703. The masks are the correct ones!!
Not a very productive night.
  Data Summary  
	
		
			|  Project  | 
			 Objects  | 
			 Filter  | 
			 ExpTime  | 
			 Notes  | 
		
	
	
		
			|  Valsecchi  | 
			 1999AO10  | 
			 g-sloan / r-sloan  | 
			 1hr  | 
			 this time NS tracking seems to have worked  | 
		
		
			|  Storm  | 
			 IC10VI  | 
			 V-bess / I-bess  | 
			 5min  | 
			 the airmass was low (~2.3) and the seeing was >2.0" so we aborted - after allseeing fwhm in blue=2.9" and red=2.6"  | 
		
		
			|  Donati  | 
			 ber27  | 
			 B / V,I  | 
			 ~10min  | 
			 after allseeing fwhm in blue=1.53" and red: V= 1.45" / I= 1.36" - considered within specs  | 
		
		
			|  Donati  | 
			 ngc2355  | 
			 B / V,I  | 
			 ~10min  | 
			 after allseeing fwhm in blue=1.65" and red: V= 1.66" / I=1.53 - marginally within specs  | 
		
		
			|  Donati  | 
			 czer24  | 
			 B / V,I  | 
			 ~10min  | 
			 after allseeing fwhm in blue=1.8" and red: V= 1.7" / I= 1.5" - out of specs  | 
		
		
			|  Donati  | 
			 ber30  | 
			 B / V,I  | 
			 ~10min  | 
			 after allseeing fwhm in blue=1.49" and red: V= 1.4" / I=1.22" - within specs  | 
		
		
			|  Donati  | 
			 coll74  | 
			 B / V,I  | 
			 ~10min  | 
			 after allseeing fwhm in blue=1.97" and red: V= 2.06" / I= 1.6" - out of specs  | 
		
		
			|  Gilli  | 
			 SDSS J  | 
			 r-sloan / i-sloan  | 
			 12min  | 
			 aborted because of bad seeing - after allseeing fwhm in blue=1.65" and red=1.4" - out of specs  | 
		
		
			|  Fiore  | 
			 Cosmos2  | 
			 dual grating  | 
			 3.5hr  | 
			 after monitoring GCS seeing measurement ~1.5hr is out of spec - see figure attached in the end of log  | 
		
		
			|  Donnarumma  | 
			 A1703m1  | 
			 red grating  | 
			 ~50min  | 
			 we aborted the second exposure because we were in twilight (only 850sec saved) - GCS seeing >1.4" during this OB - out of specs  | 
		
		
			|  STD_MODS  | 
			 Feige34  | 
			 dual/red grating  | 
		
	
  Details 
00:30  10Bias_Bino 
01:32 Flat 
-  SkyFlat_rz_pa0 scale=0.5 counts: blue 29525-14278, red 26790-13178
  -  SkyFlat_rz_pa180 scale=1.0 (maybe better scale=1.5??) counts: blue 15552-7651, red 13674-7414
  -  SkyFlat_BI_pa180 scale=6.0 counts: blue 30239 15459, red 39920-20812
  -  SkyFlat BI_pa0 scale=12.0 counts: blue 19650 10485, red 26347-14757
 
 
01:52  dofpia,/X2,/First 
 
10730      0      0      0      0      0      0   1.52   4.23     red
14012      0      0      0      0      0      0   1.77   5.52    blue
1021    -80   -318  -1687    313     13     63   1.33   0.88     red
691   -399   -158  -1397   -387     11     71   1.50   0.72    blue
-2547      7     94   -143   -257    158     86   1.22   1.02     red
-1662   -240   -123     36     34     34     80   1.61   0.67    blue
721     96     90   -135     -8    161     52   1.47   0.32     red
1125   -201    -89   -185     41     20     62   1.68   0.46    blue
-596   -103   -107    176     84     93     49   1.37   0.27     red
-684   -140   -148    194     62    -94     32   1.52   0.31    blue
-101    181     56   -145      5     53     32   1.38   0.13     red
137   -156    -67   -145   -171    -43     32   1.59   0.15    blue
  
02:05 lbcrangebal
02:14 Slew to Valsecchi: 201202122E
02:53 dofpia,/X2 on the west field
-4123     43    121    244    144   -101     -5   1.31   1.63     red
 -4434    -90    -84     -5    118    -57      0   1.21   1.75    blue
 -306     94     34    -62   -148    -35    -34   1.18   0.17     red
  171    -11    -77     38     16    -97    -36   1.43   0.12    blue
  
02:58 Valsecchi: 201202122W
03:35 Slew to 
IC10VI for dofpia
 -262    117     90    -68    -75    242     57   1.85   0.21     red
 -4260    -65     70   -102    190    367     90   1.99   1.70    blue
-1311   -112    -32    144    100    211     40   1.62   0.55     red
 -1707   -119    -33      8      2    326     60   1.83   0.71    blue
 -514    -77     93     61     50    200     58   1.81   0.26     red
 -996   -115      5   -176    -87    304     70   2.07   0.45    blue
 -856    -52     88     61    -45    227     58   1.87   0.38     red
-1009   -187    -39    261    -23    240     63   1.98   0.45    blue
  
-484    203     37    -57    -71    197     23   1.61   0.25     red
 -654    284    -89     27   -220    172     42   1.72   0.32    blue
 -1277    -77     64      5     40     90     22   1.51   0.51     red
-1055   -113    108      5    108     76     14   1.58   0.43    blue
  9    -27    141    -35    -92     65     20   1.52   0.11     red
 -379   -140    -58     75    -30     92     25   1.65   0.19    blue
03:50 IC10VI  
-  airmass ~2.3
  -  seeing on images > 2"
  -  Aborted
 
 
04:02 Slew to Donati
04:05 dofpia,/X2
5004      2   -118     -5     68   -803   -100   1.11   2.04     red
9502      0      0      0      0      0      0   1.18   3.74    blue
840    -33   -100     93    -70   -431   -100   0.99   0.44     red
2842    175      1    237   -152  -1000   -100   0.94   1.29    blue
-235     47    -93   -116     11    -92    -87   1.01   0.15     red
-549    240    -28   -351     61   -990   -100   0.96   0.69    blue
282    -53    -34    142     -9     13    -63   1.00   0.15     red
-1012    149    -71    -53    -24    -63   -100   1.21   0.41    blue
 
-  Aborted. Not conveged
  -  Try again without /X2
 
 
473    -48   -115    -11    -57    -30    -36   1.05   0.21     red
32     65    -85     42     27    -14    -92   1.29   0.09    blue
-207     90     63     35     47     23    -29   0.98   0.12     red
-233    149     42      8      0     20    -75   1.24   0.13    blue
4:16 lbcrangebal
4:18 be27  
-  seeing on images blue ~1.4 red ~1.25
 
 
04:34 slew to ngc2355_focus_ok  
 
1035    -28    -43     62    -30     -4     -6   1.19   0.42     red
1404    -78     98     71    -26    105    -20   1.42   0.56    blue
 159      0    -50   -196    131     11     -6   1.21   0.15     red
 840    138     88     41    -82     80     -3   1.43   0.35    blue
 -358     78     10    267   -102    -52     -7   1.03   0.21     red
-1268    109    -15    -64    122     -5    -15   1.25   0.51    blue
  143    -54    -33    -75    -20    -48     -4   1.15   0.11     red
 759     99    -37     88   -129     17      2   1.45   0.32    blue
  475    -28    -32     -2     57    -32    -11   1.04   0.21     red
 341   -125     62   -119    114   -120    -19   1.32   0.19    blue
  
04:45 ngc2355  
-  seeing on images ~1.6"
  -  seeing on images blue ~1.48 red ~1.35
  -  seeing on images blue ~1.6 red ~1.2
 
 
05:03 slew to czer24_focus_ok  
 
-82    116    -12     87    -70     42     10   1.16   0.11     red
702   -131     14    230     10     -2    -10   1.40   0.31    blue
138    108    100    -30    -12     86     11   1.19   0.12     red
226   -131     16    136     65     41     20   1.50   0.14    blue
  
05:09 slew to czer24 
-  seeing on images ~1.6"
  -  seeing on images blue ~1.6", red ~1.4"
  -  seeing on images blue ~1.4", red ~1.2"
 
 
05:27 slew to be30_focus_ok  
 
 
-312    -47    -34    -45     68     60     33   1.35   0.16     red
-773     18   -103   -168    148    135     43   1.62   0.34    blue
-351    -22     68     37    -53    -28     12   1.26   0.16     red
-340    -27     77    -39      7     95     26   1.59   0.17    blue
 
  
05:34 be30 
-  seeing on images blue ~1.5", red ~1.4"
  -  dimm seeing 1.6"
  -  seeing on images ~1.35"
 
 
05:49 Slew to coll74_focus_ok 
 
 
 -563    -42    -94   -111   -143    124     11   1.32   0.26     red
-1414    -74    -66   -164     89    184     22   1.56   0.58    blue
896    -42     13    -19     75    -16      1   1.20   0.36     red
-28    -34   -130    -34    -22     62     12   1.53   0.10    blue
-925      0     73    -51     29     47     15   1.31   0.38     red
-485    -31   -116    123     12     51     13   1.51   0.22    blue
-417    -72    -58    -63    142    -19     -7   1.43   0.20    blue
-66    129    -90    -61    112    -35     -7   1.20   0.11     red
-110    -60     56     55    -55    -50     -2   1.51   0.10    blue
  
06:01 coll74 
-  dimm seeing 1.8"
  -  seeing on images blue ~2", red ~1.8"
  -  seeing on images blue ~1.8", red ~1.5"
  -  arirmass on last image 1.6 
 
 
06:16 Slew to Gilli 
 
 1401     13     69    368   -228   -239    -49   1.00   0.61     red
3180    -21    405    -75   -618   -342    -68   1.29   1.31    blue
-16     24   -154   -299   -133    -39     -1   1.16   0.17     red
-522   -181    177    274    133   -102     -8   1.36   0.28    blue
-141     39     68    299     94    -55     -8   1.10   0.18     red
-259    164     93     13    -76    -55    -17   1.36   0.15    blue
  
06:28 slew to ri3-1 
-  seeing measured on images: blue:1.4 and red:1.2 
 
 
06:44 Aborted! Seeing too high
06:44 Switch to MODS
07:07 acqMODS mods.57.FIORE2.acq 
mods1r 0002-0004 
-  MODS offset: -0.2211  0.316 -0.967 detxy rel
 
 
07:20 execMODS mods.57.FIORE2.obs 
mods1b 0002-0008 / mods1r 0005-0011 
-  Last exposure aborted. Airmass limit reached
 
 
11:04 acqMODS Feige34_std.acq 
mods1r 0012-0017
11:19 execMODS Feige34_std.obs 
mods1b 0009 / mods1r 0018 
-  airmass 1.35
  -  the red spectrum in mods1r_0018 is with dual_grating for Fiore 
 
 
11:27 execMODS Feige34_stdRED.obs 
mods1r 0019 
-  airmass 1.35
  -  mods1r_0019 is with red_grating for Donnarumma but it is saturated 
 
 
11:34 divided exp time of Feige34_stdRED.obs by 2 
mods1r 0020
11:40 acqMODS mods.201151.newmi1e.acq 
mods1r 0021-0023
11:55 execMODS mods.201151.newmi1e.obs 
mods1r 0024-0025
12:51 Abort OB because of twilight. The second exposure is only 856sec instead of 2250se (mods1r0025)
01:07 Close Dome - Starting calibrations 
-  The Donnarumma calibrations for mask1 have as PROPID sdss1029 even though the actual target is A1703. The maskid is the correct one!!
  -  execMODS -u grmosflatsDonnaRm1.cal 
mods1r 0026-0035
  -  execMODS -u grmoslampsDonnaRm1.cal 
mods1r 0036-0038
  -  execMODS -u grpixflats.cal 
mods1b 0010-0019 / mods1r 0039-0043
  -  execMODS -u allbias.cal 
mods1b 0020-0039 / mods1r 0044-0063
 
 
02:22 End of calibrations - put MODS to sleep 
-  execMODS ~MODSeng/modsScripts/Support/modsSleep.pro 
 
 
 
-  Seeing reported by GCS while we were obtaining FIORE2: 
     
 
 
 
-  Seeing reported by GCS while we were obtaining Donnarumma: 
     
 
 
-- 
KonstantinaBoutsia - 22 Feb 2012